OpenStudy (anonymous):

Describe how the graph of y=|x| – 7 is like the graph of y=|x| and how it is different.

OpenStudy (anonymous):

down by 7 units

OpenStudy (anonymous):

You asked the very same question a few moments ago. So when you translate the origin to a different point, the graph stays the same, but instead starting in the origin (0,0) it will start in (0,-7)

OpenStudy (anonymous):

so @Diogo the fact you said i asked the similar question is it the same anwser ?

OpenStudy (kainui):

@mathisblockingmypath you should try to figure these out on your own by graphing them. What happens when you plug in x=0 to this equation? Do you even know? You're just floating through life like it doesn't matter, like you're helpless. You're not.

OpenStudy (anonymous):

i wanna know if the x or y inteercept would change

OpenStudy (anonymous):

@Kainui i used this site to make sure im right or wrong thank you

OpenStudy (mathmale):

"Describe how the graph of y=|x| – 7 is like the graph of y=|x| and how it is different." Your topic here is "horizontal and vertical translations." If you begin with y=f(x) and graph it, and then replace (x) with (x-a), your original graph will need to be translated (moved) 'a' units to the right. That's not what's happening in your question. If you begin with y=f(x) and then modify this function by adding 'a' to it, obtaining y=f(x)+a, then your graph will need to be translated (moved) vertically upward by 'a' units. You might want to write all this down, as you will surely need it in the future. Try again: Given y=|x|, graph it. Now consider how you would have to move / translate this graph up or down, left or right, to represent y=|x|-7.